A 25.0 mL sample of 0.150 M HIO3 is titrated with 0.240 M Sr(OH)2. What volume of Sr(OH)2 is required to reach the equivalence point?
1
Expert's answer
2020-05-28T10:31:52-0400

"2HIO_3+Sr(OH)_2\to Sr(IO_3)_2+2H_2O"

Let "M_1=0.15M,V_1=25\ ml,n_1=2;"

"M_2=0.24 M,n_2=1,V_2=?"

By Molarity equation,

"\frac{M_1V_1}{n_1}=\frac{M_2V_2}{n_2}\implies"

"\frac{0.15\times 25}{2}=\frac{0.24\times V_2}{1}" "\implies V_2=7.81 \ ml"